OpenStudy (joshoyen):

Find the zeros. 1. y=x^4-8x^2+16

sam (.sam.):

y=x^4-8x^2+16 \[ x^4-8x^2+16=0\] Now factorize it

OpenStudy (joshoyen):

@.Sam. Yeah, that's where I'm having trouble at.

sam (.sam.):

\[(x^2-4)(x^2-4)=0\]?

sam (.sam.):

looks good?

OpenStudy (joshoyen):

ohh okay, yeah thanks man

sam (.sam.):

No problem